OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 5 inches and the radius is 6 inches. If none of the photos overlap, what is the area of photographs you need? Use for 3.14 for pi

jimthompson5910 (jim_thompson5910):

what you want is the lateral surface area

jimthompson5910 (jim_thompson5910):

the lateral surface area for a cylinder is LSA = 2*pi*r*h in this case, r = 6 h = 5

jimthompson5910 (jim_thompson5910):

so LSA = 2*pi*r*h LSA = 2*pi*6*5 LSA = 60*pi LSA = 60*3.14 LSA = 188.4 So the lateral surface area is roughly 188.4 square inches